OpenStudy (anonymous):

Choose the proper expression for the nth term (recall, ^ means raised to): 16, 64, 256

OpenStudy (unklerhaukus):

can you express those terms as powers of some number?

OpenStudy (unklerhaukus):

16 = a^n 64 = a^(n+1) 256 = a^(n+2)

OpenStudy (anonymous):

these are the choices they give